At standard atmospheric pressure, saturated steam has a specific volume of 26.80 ft3 /lb. If the enthalpy of that same vapor is 1150.5 Btu/lb, calculate the internal energy of the steam in Btu/lb.
1
Expert's answer
2020-03-17T09:51:34-0400

"26.80 \frac{ft^3}{lb}=1.673 \frac{m^3}{kg}"


"1150.5 \frac{Btu}{lb}=2676.063 \frac{kJ}{kg}"


So,


"H=U+pV\to \frac{H}{m}=\frac{U}{m}+p\frac{V}{m}\to h=u+pv"


"u=h-pv=2676.063\cdot10^3-101325\cdot 1.673=2506.546 \frac{kJ}{kg}"



"2506.546 \frac{kJ}{kg}=1077.62" "\frac{Btu}{lb}"


"u=1077.62" "\frac{Btu}{lb}" Answer